OpenStudy (anonymous):

room temperature ranges from 20 degrees celsius to 25 degrees celsius. Find the range of room temperature in Fahrenheit. Use the formula, F-32=1.8C, to convert from the Celsius scale to the Fahrenheit scale.

OpenStudy (anonymous):

just plug in 20 and 25 for C and solve for F. Those F values will define the range

OpenStudy (uh60blackhawk):

[20x5/9]=36 36+32=68 Fahrenheit now try the last one by yourself.

OpenStudy (uh60blackhawk):

The formula to convert Celsius to Fahrenheit is C x 9/5 + 32 = °F

OpenStudy (uh60blackhawk):

You there?

OpenStudy (anonymous):

Got it...thanks. 68- 77 degrees F.

OpenStudy (uh60blackhawk):

your welcome.